OpenStudy (anonymous):

solve for f 4f/5>8 f = 10 f > 10 f < 10 None of the above

OpenStudy (anonymous):

F can't equal 10 since that would end up being 8=8 so F must be larger than 10

OpenStudy (anonymous):

ok so start out replacing your> with a =

OpenStudy (anonymous):

4f/5=8

OpenStudy (anonymous):

I get f=10 is that the answer @bannanasplitz

OpenStudy (anonymous):

No, F can't be equal to 10 F is larger than 10

OpenStudy (anonymous):

try doing the equation 4Xf/5=8 is that correct?

OpenStudy (anonymous):

correction is 4x10/5=8

OpenStudy (anonymous):

so it actually is correct good job @undeadking1357

OpenStudy (anonymous):

But it says it must be larger than =8 So F is bigger than 10 (f > 10) You can't change 4f/5>8 to 4f/5=8. It changes the whole problem

OpenStudy (anonymous):

F does not equal 10

OpenStudy (anonymous):

so @silenceforthedead the answer would be f<10?

OpenStudy (anonymous):

Nope your answer would be f > 10

OpenStudy (anonymous):

Oh thank you @silenceforthedead it was right!

OpenStudy (anonymous):

You're welcome ^-^